# Modelling Non-monotone Risk Aversion and Convex Compensation in Incomplete Markets

### v1

@@ 3.2 PSAHARA Utility
A function $U:\mathbb{R}\rightarrow\mathbb{R}$ is a piecewise SAHARA utiltiy if and only if there exists a partition $\{a_{k}\}_{k=0}^{n+1}$ and a family of parameter tuples $\{\left(\alpha_{k},\beta_{k},d_{k},u_{k}\right)\}_{k=0}^{n}$ such that
@@
The concave envelope plays an important role in the portfolio choice problem with non-concave utilities. As shown in Appendix A, if $\xi_{T}$ has a continuous distribution, Problem (16) has the same optimal solution as the following problem with the utility $U$ replaced by the concave envelope $U^{**}$:
@@
This example further shows the generality of the PSAHARA utility class. We consider a non-monotone and discontinuous utility function:
@@
where we let $l_{1}=-6,l_{2}=-4.5,l_{3}=-1,l_{4}=2,\beta=1,\alpha_{0}=1.7,\alpha_{3}=2.2,\alpha_{5}=1.2,d_{0}=3,d_{3}=1,d_{5}=6,\gamma_{0}=2,\gamma_{3}=1.5,\gamma_{5}=7$. $b_{1},b_{2},b_{3},b_{4}$ are arbitrary real numbers to make the function continuous on $x\geqslant l_{1}$. This utility function is not a PSAHARA utility since there exist decreasing segments and jumps. However, as shown in Figure 2, the concave envelope of (20) is a PSAHARA utility.
$\tilde{U}\left(x\right):=\left\{\begin{aligned} &\gamma_{0}\hat{U}\left(x;\alpha_{0},\beta,d_{0}\right)+u_{0},&&x<a_{1};\\ &\gamma_{0}\hat{U}^{\prime}\left(a_{1};\alpha_{0},\beta,d_{0}\right)\left(x-a_{1}\right)\\ &\quad+\hat{U}\left(a_{1};\alpha_{0},\beta,d_{0}\right),&&a_{1}\leqslant x<a_{2};\\ &\gamma_{3}\hat{U}^{\prime}\left(a_{3};\alpha_{3},\beta,d_{3}\right)\left(x-a_{3}\right)\\ &\quad+\hat{U}\left(a_{3};\alpha_{3},\beta,d_{3}\right),&&a_{2}\leqslant x<a_{3};\\ &\gamma_{3}\hat{U}\left(x;\alpha_{3},\beta,d_{3}\right)+u_{3},&&a_{3}\leqslant x<a_{4};\\ &\gamma_{5}\hat{U}^{\prime}\left(a_{5};\alpha_{5},\beta,d_{5}\right)\left(x-a_{5}\right)\\ &\quad+\hat{U}\left(a_{5};\alpha_{5},\beta,d_{5}\right),&&a_{4}\leqslant x<a_{5};\\ &\gamma_{5}\hat{U}\left(x;\alpha_{5},\beta,d_{5}\right)+u_{5},&&x\geqslant a_{5}.\end{aligned}\right.$
Figure 2: The original utility (20) and the concave envelope. In the figure, the red parts are the segments where the concave envelope coincides with the original utility, while the blue straight lines are the segments where the concave envelope does not coincide with the original utility. $\hat{U}$ is the utility function defined in (2) and $\hat{U}^{\prime}$ is the derivative of $\hat{U}$ shown in (3). $a_{1},a_{3}$, $a_{5}$ are the tangent points between the linear segments and the strictly concave parts. $(l_{1},l_{3})=(a_{2},a_{4})$. $u_{0},u_{3},u_{5}$ are the correction constants to make the function continuous. As labeled in the figure, $U$ is non-differentiable at $a_{2}$ and $a_{4}$.
@@ 4 A Unified Formula of the Optimal Portfolio
We proceed to show the optimal terminal wealth, the optimal wealth process and the optimal control of PSAHARA utilities for Problem (16). We adopt the martingale and duality method and the concavification technique. The proofs are included in Appendix A.
@@
  • The terms $X_{t,k}^{R}$ and $X_{t,k}^{\bar{R}}$ arise from the symmetric nature of SAHARA utility functions, representing the investor’s propensity towards the risk-seeking behavior. Specifically, $X_{t,k}^{R}$ encourages an increase to the risky investment when the wealth exceeds the threshold level, and hence is called the positive main term. Conversely, $X_{t,k}^{\bar{R}}$ promotes an increase to the risky investment when wealth is below the threshold, hence referred to as the negative main term. These two terms are the main driver of $X_{t}^{*}$, as illustrated in the later Section 5 of asymptotic analysis.
@@
In the following theorem, we show the optimal control of Problem (19), which is also the unified formula of the optimal portfolio for PSAHARA utilities. The proof is included in Appendix A.
@@
where $b_{t}:=\beta e^{-\int_{t}^{T}\left(r_{s}-\frac{\|\theta_{s}\|_{2}^{2}}{\alpha^{2}}\right)\,\mathrm{d}s}$. This coincides with Theorem 3.2 in Chen, Pelsser and Vellekoop (2011).
@@
We name the term $\boldsymbol{\pi}_{t}^{(1)}$ the asymptotic Merton term. As illustrated later in Section 5, this term drives $\boldsymbol{\pi}_{t}^{*}/X_{t}^{*}$ to the deterministic proportion $\frac{\boldsymbol{\sigma}_{t}^{\intercal}\left(\boldsymbol{\sigma}_{t}\boldsymbol{\sigma}_{t}^{\intercal}\right)^{-1}\left(\boldsymbol{\mu}_{t}-r_{t}\mathbf{1}_{m}\right)}{\alpha_{n}}$ as $\xi_{t}$ approaches 00 and $-\frac{\boldsymbol{\sigma}_{t}^{\intercal}\left(\boldsymbol{\sigma}_{t}\boldsymbol{\sigma}_{t}^{\intercal}\right)^{-1}\left(\boldsymbol{\mu}_{t}-r_{t}\mathbf{1}_{m}\right)}{\alpha_{0}}$ as $\xi_{t}$ approaches $\infty$, corresponding to the famous Merton proportion (Merton (1969)). We claim that $\boldsymbol{\pi}_{t}^{(1)}$ serves as the main driver behind the dynamics of the optimal portfolio.
The second term $\boldsymbol{\pi}_{t}^{(2)}$ is identified as the risk adjustment term. It leads to risk-seeking behaviors at high wealth levels and results in risk-aversion at low wealth levels. Containing the scale parameter $\beta$ in $X_{t,k}^{R}$, this term shows a distinct feature that it adjusts the optimal portfolio corresponding to the total wealth scale. This term is the second most important driver in the PSAHARA portfolio.
@@
5 Asymptotic Analysis

### v2

@@ 3.2 PSAHARA Utility
A function $U:\mathbb{R}\rightarrow\mathbb{R}$ is a piecewise SAHARA utilitiy if and only if there exists a partition $\{a_{k}\}_{k=0}^{n+1}$ and a family of parameter tuples $\{\left(\alpha_{k},\beta_{k},d_{k},u_{k}\right)\}_{k=0}^{n}$ such that
@@
The concave envelope plays an important role in the portfolio choice problem with non-concave utilities. We rigorously show in the proof of Theorem 1 that if $\xi_{T}$ has a continuous distribution, Problem (22) has the same optimal solution as the following problem with the utility $U$ replaced by the concave envelope $U^{**}$:
@@
This example further shows the generality of the PSAHARA utility family. We consider a non-monotone and discontinuous utility function:
@@
where we let $l_{1}=-6,l_{2}=-4.5,l_{3}=-1,l_{4}=2,\beta=1,\alpha_{0}=1.7,\alpha_{3}=2.2,\alpha_{5}=1.2,d_{0}=3,d_{3}=1,d_{5}=6,\gamma_{0}=2,\gamma_{3}=1.5,\gamma_{5}=7$. $b_{1},b_{2},b_{3},b_{4}$ are suitable real numbers to make the function continuous on $x\geqslant l_{1}$. This utility function is not a PSAHARA utility since there exist decreasing segments and jumps. However, as shown in Figure 2, the concave envelope of (26) is a PSAHARA utility.
$U^{**}\left(x\right):=\left\{\begin{aligned} &\gamma_{0}\hat{U}\left(x;\alpha_{0},\beta,d_{0}\right)+u_{0},&&x<a_{1};\\ &\gamma_{0}\hat{U}^{\prime}\left(a_{1};\alpha_{0},\beta,d_{0}\right)\left(x-a_{1}\right)\\ &\quad+\hat{U}\left(a_{1};\alpha_{0},\beta,d_{0}\right),&&a_{1}\leqslant x<a_{2};\\ &\gamma_{3}\hat{U}^{\prime}\left(a_{3};\alpha_{3},\beta,d_{3}\right)\left(x-a_{3}\right)\\ &\quad+\hat{U}\left(a_{3};\alpha_{3},\beta,d_{3}\right),&&a_{2}\leqslant x<a_{3};\\ &\gamma_{3}\hat{U}\left(x;\alpha_{3},\beta,d_{3}\right)+u_{3},&&a_{3}\leqslant x<a_{4};\\ &\gamma_{5}\hat{U}^{\prime}\left(a_{5};\alpha_{5},\beta,d_{5}\right)\left(x-a_{5}\right)\\ &\quad+\hat{U}\left(a_{5};\alpha_{5},\beta,d_{5}\right),&&a_{4}\leqslant x<a_{5};\\ &\gamma_{5}\hat{U}\left(x;\alpha_{5},\beta,d_{5}\right)+u_{5},&&x\geqslant a_{5}.\end{aligned}\right.$
Figure 2: The original utility (26) and the concave envelope. In the figure, the red parts are the segments where the concave envelope coincides with the original utility, while the blue straight lines are the segments where the concave envelope does not coincide with the original utility. $\hat{U}$ is the utility function defined in (2) and $\hat{U}^{\prime}$ is the derivative of $\hat{U}$ shown in (3). $a_{1},a_{3}$, $a_{5}$ are the tangent points between the linear segments and the strictly concave parts. $(l_{1},l_{3})=(a_{2},a_{4})$. $u_{0},u_{3},u_{5}$ are the correction constants to make the function continuous. As labeled in the figure, $U^{**}$ is non-differentiable at $a_{2}$ and $a_{4}$.
@@ 4 A Unified Formula of the Optimal Portfolio
4.1 Optimal Wealth and Optimal Portfolio
We proceed to show the optimal terminal wealth, the optimal wealth process and the optimal control of PSAHARA utilities for Problem (22). We adopt the martingale and duality method and the concavification technique. The proofs are included in Section 4.2 and Appendix A.
@@ 4.1 Optimal Wealth and Optimal Portfolio
  • The terms $X_{t,k}^{R}$ and $X_{t,k}^{\bar{R}}$ arise from the symmetric nature of SAHARA utility functions, representing the investor’s propensity towards the risk-seeking behavior. Specifically, $X_{t,k}^{R}$ encourages an increase to the risky investment when the wealth exceeds the threshold level, and hence is called the positive main term. Conversely, $X_{t,k}^{\bar{R}}$ promotes an increase to the risky investment when wealth is below the threshold, and it is hence referred to as the negative main term. These two terms are the main driver of $X_{t}^{*}$, as illustrated in the later Section 5.1 of asymptotic analysis.
@@
In the following theorem, we show the optimal control of Problem (25), which is also the unified formula of the optimal portfolio for PSAHARA utilities. The proof is included in Section 4.2 and Appendix A.
@@
where $b_{t}:=\beta e^{-\int_{t}^{T}\left(r_{s}-\frac{\|\boldsymbol{\theta}_{s}\|_{2}^{2}}{\alpha^{2}}\right)\,\mathrm{d}s}$. This coincides with Theorem 3.2 in Chen, Pelsser and Vellekoop (2011).
@@
We name the term $\boldsymbol{\pi}_{t}^{(1)}$ the asymptotic Merton term. As illustrated later in Section 5.1, this term drives $\boldsymbol{\pi}_{t}^{*}/X_{t}^{*}$ to the deterministic proportion $\frac{\boldsymbol{\sigma}_{t}^{\intercal}\left(\boldsymbol{\sigma}_{t}\boldsymbol{\sigma}_{t}^{\intercal}\right)^{-1}\left(\boldsymbol{\mu}_{t}-r_{t}\mathbf{1}_{m}\right)}{\alpha_{n}}$ as $\xi_{t}$ approaches 00 and $-\frac{\boldsymbol{\sigma}_{t}^{\intercal}\left(\boldsymbol{\sigma}_{t}\boldsymbol{\sigma}_{t}^{\intercal}\right)^{-1}\left(\boldsymbol{\mu}_{t}-r_{t}\mathbf{1}_{m}\right)}{\alpha_{0}}$ as $\xi_{t}$ approaches $\infty$, corresponding to the famous Merton proportion (Merton (1969)). We claim that $\boldsymbol{\pi}_{t}^{(1)}$ serves as the main driver behind the dynamics of the optimal portfolio.
The second term $\boldsymbol{\pi}_{t}^{(2)}$ is identified as the risk adjustment term. It leads to risk-seeking behaviors at high wealth levels and results in risk-aversion at low wealth levels. Containing the scale parameter $\beta$ in $X_{t,k}^{R}$, this term shows a distinct feature that it adjusts the optimal portfolio corresponding to the total wealth scale.
@@
4.2 Technical Discussions
This subsection discusses technical details about how Assumptions 2 and 3 work in proving Theorems 1-2. Before presenting the proof, we first claim that one difficulty to show Theorem 1 is the market incompleteness, which leads to non-unique pricing kernels and the fact that not every contingent claim can be replicated by a self-financing trading strategy. In the following, we present the proof in successive steps. First, we show that the product of any pricing kernel and any admissible wealth process, $\{\zeta_{t}X_{t}\}_{0\leqslant t\leqslant T}$, is a supermartingale. Second, we find the optimal terminal wealth for the concave envelope and show that it also solves the optimization problem of the original utility. In other words, we show that the concavification principle (see Carpenter (2000)) holds in incomplete markets. Last, we show that this terminal wealth can be replicated. Here we present the proof of supermartingality to stress the importance of Assumptions 2-3. The remaining part of the proof of Theorem 1 is in Appendix A.3.
We begin with an elementary proposition. Let $\mathcal{M}$ denote the set of all pricing kernel processes. That is,
[…]
Let $\{X_{t}^{x_{0}}\}_{0\leqslant t\leqslant T}$ denote any wealth process with initial value $x_{0}$. Let $\mathcal{X}^{x_{0}}$ denote the set of all terminal wealth variables, i.e.,
$\mathcal{X}^{x_{0}}:=\left\{X_{T}^{x_{0}}=x_{0}+\int_{0}^{T}\left(r_{t}X_{t}^{x_{0}}+\boldsymbol{\pi}_{t}^{\intercal}\left(\boldsymbol{\mu}_{t}-r_{t}\mathbf{1}_{m}\right)\right)\mathrm{d}t+\int_{0}^{T}\boldsymbol{\pi}_{t}^{\intercal}\boldsymbol{\sigma}_{t}\mathrm{d}\mathbf{W}_{t}:\boldsymbol{\pi}\in\mathcal{V}\right\}.$ (38)
Proposition 3.
For any $\{\zeta_{t}\}_{0\leqslant t\leqslant T}\in\mathcal{M}$ and $X_{T}^{x_{0}}\in\mathcal{X}^{x_{0}}$, the process $\{\zeta_{t}X_{t}^{x_{0}}\}_{0\leqslant t\leqslant T}$ is a supermartingale under either Assumption 2 or Assumption 3.
Proof.
Applying Itô’s formula, we get
$\mathrm{d}\left(\zeta_{t}X_{t}^{x_{0}}\right)=\left[\zeta_{t}\boldsymbol{\pi}_{t}^{\intercal}\boldsymbol{\sigma}_{t}-\zeta_{t}X_{t}^{x_{0}}\tilde{\boldsymbol{\theta}}_{t}^{\intercal}\right]\mathrm{d}\mathbf{W}_{t},$ (39)
which shows that $\{\zeta_{t}X_{t}^{x_{0}}\}_{0\leqslant t\leqslant T}$ is a local martingale. If Assumption 3 holds, then $\{\zeta_{t}X_{t}^{x_{0}}\}_{0\leqslant t\leqslant T}$ is a local martingale with a lower bound and hence a supermartingale. This is a common assumption as the utility in the literature usually has a domain bounded from left. As we allow the utility functions to have a unbounded domain from left (i.e., $X_{t}^{x_{0}}$ can be arbitrarily negative and there is no lower bound), we need to propose some other assumption to make $\{\zeta_{t}X_{t}^{x_{0}}\}_{0\leqslant t\leqslant T}$ in Eq. (39) indeed a martingale. Hence, Assumption 2 is proposed from the perspective of integrability. For the rest of this proof, we prove the statement under Assumption 2. Note that
[…]
By Hölder’s inequality, for any $p,q>1$ with $\frac{1}{p}+\frac{1}{q}=1$, we have
[…]
Moreover, we have
$\mathbb{E}\left[\int_{0}^{T}\zeta_{t}^{2q}\,\mathrm{d}t\right]$ $=\mathbb{E}\left[\int_{0}^{T}\exp\left\{-q\int_{0}^{t}\|\tilde{\boldsymbol{\theta}}_{s}\|_{2}^{2}\,\mathrm{d}s-2q\int_{0}^{t}\tilde{\boldsymbol{\theta}}_{s}^{\intercal}\mathrm{d}\mathbf{W}_{s}\right\}\,\mathrm{d}t\right]$
$\leqslant\mathbb{E}\left[\left(\sup_{t\in[0,T]}\exp\left\{-q\int_{0}^{t}\|\tilde{\boldsymbol{\theta}}_{s}\|_{2}^{2}\,\mathrm{d}s\right\}\right)\int_{0}^{T}\exp\left\{-2q\int_{0}^{t}\tilde{\boldsymbol{\theta}}_{s}^{\intercal}\,\mathrm{d}\mathbf{W}_{s}\right\}\,\mathrm{d}t\right]$
$\leqslant\mathbb{E}\left[1\cdot\int_{0}^{T}\exp\left\{-2q\int_{0}^{t}\tilde{\boldsymbol{\theta}}_{s}^{\intercal}\,\mathrm{d}\mathbf{W}_{s}\right\}\,\mathrm{d}t\right]$
[…]
Let $\rho_{t}:=\int_{0}^{t}\tilde{\boldsymbol{\theta}}_{s}^{\intercal}\,\mathrm{d}\mathbf{W}_{s}$. Then $\rho_{t}\sim\mathcal{N}\left(0,\int_{0}^{t}\|\tilde{\boldsymbol{\theta}}_{s}\|_{2}^{2}\,\mathrm{d}s\right)$. Let $\tilde{\sigma}_{t}:=\sqrt{\int_{0}^{t}\|\tilde{\boldsymbol{\theta}}_{s}\|_{2}^{2}\,\mathrm{d}s}$. By direct computation, we have
$\mathbb{E}\left[\exp\left\{-2q\rho_{t}\right\}\right]$ $=\int_{\mathbb{R}}\frac{1}{\sqrt{2\pi\tilde{\sigma}_{t}^{2}}}\exp\left\{-2qx-\frac{x^{2}}{2\tilde{\sigma}_{t}^{2}}\right\}\,\mathrm{d}x=\exp\left\{2\tilde{\sigma}_{t}^{2}q^{2}\right\}.$ (43)
It follows that
[…]
Now, we claim that the finiteness of the two terms on the right hand side of (40) relies on that $\mathbb{E}\left[\int_{0}^{T}\|\boldsymbol{\pi}_{t}\|_{2}^{2p}\,\mathrm{d}t\right]<\infty$ for some $p>1$; we will prove this later in details. Hence, we propose Assumption 2. Under this assumption, we can find some $p=1+\frac{\varepsilon}{2}$ and a corresponding $q$ such that every term on the right hand side of Eq. (41) is finite. On the other hand, similar to Eq. (41), we have
[…]
Solving (18) gives
$X_{t}^{x_{0}}=e^{\int_{0}^{t}r_{s}\,\mathrm{d}s}\left[x_{0}+\int_{0}^{t}e^{-\int_{0}^{s}r_{u}\,\mathrm{d}u}\left(\boldsymbol{\mu}_{s}-r_{s}\mathbf{1}_{m}\right)^{\intercal}\boldsymbol{\pi}_{s}\,\mathrm{d}s+\int_{0}^{t}e^{-\int_{0}^{s}r_{u}\,\mathrm{d}u}\boldsymbol{\pi}_{s}^{\intercal}\boldsymbol{\sigma}_{s}\mathrm{d}\mathbf{W}_{s}\right].$ (46)
With this, the following inequality holds under Assumption 2:
$\int_{0}^{T}\mathbb{E}\left[\left(X_{t}^{x_{0}}\right)^{2p}\right]\,\mathrm{d}t$ $\leqslant T\cdot\mathbb{E}\left[\left(\sup_{t\in[0,T]}|X_{t}^{x_{0}}|\right)^{2p}\right]$ (47)
[…]
$=T\cdot\tilde{K}_{p}\cdot\mathbb{E}\left[\left(\int_{0}^{T}\|\boldsymbol{\pi}_{t}^{\intercal}\boldsymbol{\sigma}_{t}\|_{2}^{2}\,\mathrm{d}t\right)^{p}\right]$
$\leqslant T\cdot\tilde{K}_{p}\cdot\mathbb{E}\left[\left(\int_{0}^{T}\lambda_{t}^{\max}\|\boldsymbol{\pi}_{t}\|_{2}^{2}\,\mathrm{d}t\right)^{p}\right]$
$=T\cdot\tilde{K}_{p}\left(\sup_{t\in[0,T]}\lambda_{t}^{\max}\right)^{p}\cdot\mathbb{E}\left[\left(\int_{0}^{T}\|\boldsymbol{\pi}_{t}\|_{2}^{2}\,\mathrm{d}t\right)^{p}\right]$
[…]
$=T^{1+\frac{p}{q}}\cdot\tilde{K}_{p}\cdot\left(\sup_{t\in[0,T]}\lambda_{t}^{\max}\right)^{p}\mathbb{E}\left[\int_{0}^{T}\|\boldsymbol{\pi}_{t}\|_{2}^{2p}\,\mathrm{d}t\right]<\infty,$
where $\tilde{K}_{p}$ is a constant depending on $p$. Again, if $\mathbb{E}\left[\int_{0}^{T}\|\boldsymbol{\pi}_{t}\|_{2}^{2+\varepsilon}\,\mathrm{d}t\right]$ is finite for some $\varepsilon>0$, we have that $\mathbb{E}\left[\int_{0}^{T}\zeta_{t}^{2}\left(X_{t}^{x_{0}}\right)^{2}\|\boldsymbol{\theta}_{t}\|_{2}^{2}\mathrm{d}t\right]$ is finite. By Corrollary 3.2.6 in Øksendal (2003), the two finiteness in Eqs. (41) and (45) guarantee that $\{\zeta_{t}X_{t}^{x_{0}}\}_{0\leqslant t\leqslant T}$ is an $\{\mathcal{F}_{t}\}_{0\leqslant t\leqslant T}$-martingale, and hence a supermartingale. ∎
Further, we give a proposition on the concavification technique in incomplete markets.
Proposition 4.
Let $U^{**}$ be the concave envelope of $U$. Then the optimization problem (25) shares the same optimal solution as problem (22).
Proof.
We first derive the concave envelope $U^{**}$ of $U$ and discuss the optimization problem (25). To derive the former, we utilize the convex conjugate of $U^{**}$:
$\tilde{U}^{**}\left(\kappa\right)$ $=\max_{x\in\text{dom}U}\left\{U^{**}\left(x\right)-\kappa x\right\}$ (48)
$=U^{**}\left(I\left(\kappa\right)\right)-\kappa I\left(\kappa\right),\quad\kappa>0,$
where $I$ is the generalized inverse of $\left(U^{**}\right)^{\prime}$, defined by
$I(y):=\inf\left\{x\in\text{dom}U:\left(U^{**}\right)^{\prime}(x)\leqslant y\right\}.$ (49)
By definition of $I(\cdot)$, the function $y\mapsto\mathbb{E}\left[\zeta_{T}I\left(y\zeta_{T}\right)\right]$ is strictly decreasing for any $\zeta\in\mathcal{M}$. Hence, for fixed $\zeta\in\mathcal{M}$, we can define a function $\mathcal{Y}_{\zeta}\left(\cdot\right)$ as the inverse of $y\mapsto\mathbb{E}\left[\zeta_{T}I\left(y\zeta_{T}\right)\right]$. Substituting $\kappa=\mathcal{Y}_{\zeta}\left(x_{0}\right)\zeta_{T}$ in (48), we have
$U^{**}\left(I\left(\mathcal{Y}_{\zeta}\left(x_{0}\right)\zeta_{T}\right)\right)-\mathcal{Y}_{\zeta}\left(x_{0}\right)\zeta_{T}I\left(\mathcal{Y}_{\zeta}\left(x_{0}\right)\zeta_{T}\right)\geqslant U^{**}\left(X_{T}^{x_{0}}\right)-\mathcal{Y}_{\zeta}\left(x_{0}\right)\zeta_{T}X_{T}^{x_{0}}.$ (50)
Additionally, $\mathbb{E}\left[\zeta_{T}I\left(\mathcal{Y}_{\zeta}\left(x_{0}\right)\zeta_{T}\right)\right]=x_{0}$ and $\mathbb{E}\left[\zeta_{T}X_{T}^{x_{0}}\right]\leqslant x_{0}$. Taking expectation on both sides of (50), we have
$\mathbb{E}\left[U^{**}\left(I\left(\mathcal{Y}_{\zeta}\left(x_{0}\right)\zeta_{T}\right)\right)\right]\geqslant\mathbb{E}\left[U^{**}\left(X_{T}^{x_{0}}\right)\right].$ (51)
Since $X_{T}^{x_{0}}\in\mathcal{X}^{x_{0}}$ is arbitrary, this proves the optimality of the terminal wealth $I\left(\mathcal{Y}_{\zeta}\left(x_{0}\right)\zeta_{T}\right)$. Moreover, by the arbitrariness of $\zeta\in\mathcal{M}$, we have that $\mathbb{E}\left[U^{**}\left(I\left(\mathcal{Y}_{\zeta}\left(x_{0}\right)\zeta_{T}\right)\right)\right]$ has the same value for all $\zeta\in\mathcal{M}$. Hence, we use $\xi$ defined in (17) as a representative. By definition, we know that $U^{**}(x)\geqslant U(x)$ for all $x\in\text{dom}U$. Hence, we have
$\mathbb{E}\left[U^{**}\left(I\left(\mathcal{Y}_{\xi}\left(x_{0}\right)\xi_{T}\right)\right)\right]\geqslant\mathbb{E}\left[U^{**}\left(X_{T}^{x_{0}}\right)\right]\geqslant\mathbb{E}\left[U\left(X_{T}^{x_{0}}\right)\right].$ (52)
Define the index set
Note that $U^{**}\left(x\right)\neq U\left(x\right)$ if and only if $x\in\bigcup_{k\in\mathcal{I}}\left(a_{k},a_{k+1}\right)$. Moreover, since $\xi_{T}$ is continuously distributed, we have
$\mathbb{P}\left(\mathcal{Y}_{\xi}(x_{0})\xi_{T}\in\{\gamma_{k}^{-},\gamma_{k}^{+}\}\right)\leqslant\mathbb{P}\left(\mathcal{Y}_{\xi}(x_{0})\xi_{T}\in\{\gamma_{k}^{-},\gamma_{k}^{+}\}\cup\{\gamma_{n+1}^{-}\}\right)=0.$ (54)
Now for notation simplicity, let $y^{*}=\mathcal{Y}_{\xi}(x_{0})$ and $X_{T}^{*}=I\left(y^{*}\xi_{T}\right)$. We have
$=\mathbb{E}\left[U^{**}\left(X_{T}^{*}\right)\right].$ (55)
This proves that the optimizer of Problem (25) also maximizes Problem (22). ∎
A seminal literature on expected utility maximization in incomplete markets is Karatzas, Lehoczky, Shreve and Xu (1991), in which the so–called fictitious completion thought experiment is introduced. The authors suppose that the market is completed by $(q-m)$ risky assets with the return vector $\boldsymbol{\alpha}_{t}\in\mathbb{R}^{q-m}$ and volatility matrix $\boldsymbol{\rho}_{t}\in\mathbb{R}^{(q-m)\times q}$, where the rows of $\boldsymbol{\rho}_{t}$ are orthonormal and in the kernel of $\boldsymbol{\sigma}_{t}$. This orthogonality, in some way, implies that the optimal portfolio in incomplete markets should coincide with the one in complete markets. We proceed to show this implication in details.
Suppose that the market is completed by $(q-m)$ risky assets with the return vector $\boldsymbol{\alpha}_{t}\in\mathbb{R}^{q-m}$ and volatility matrix $\boldsymbol{\rho}_{t}\in\mathbb{R}^{(q-m)\times q}$, where the rows of $\boldsymbol{\rho}_{t}$ are orthonormal and in the kernel of $\boldsymbol{\sigma}_{t}$. Let
$\boldsymbol{\tilde{\theta}}_{\boldsymbol{\nu},t}:=\boldsymbol{\tilde{\sigma}}_{t}^{\intercal}\left(\boldsymbol{\tilde{\sigma}}_{t}\boldsymbol{\tilde{\sigma}}_{t}^{\intercal}\right)^{-1}\left(\boldsymbol{\tilde{\mu}}_{t}-r_{t}\mathbf{1}_{q}\right)\\ =\boldsymbol{\theta}_{t}+\boldsymbol{\nu}_{t},$ (57)
Lemma 1.
Proposition 5.
Under the assumption given by Eq. (15), then $\{\boldsymbol{\pi}_{t}^{*}\}_{0\leqslant t\leqslant T}$ in Theorem 2 is the optimal portfolio in the incomplete market, i.e.,
Proof.
We only need to show that if $\{\boldsymbol{\nu}_{t}\}_{0\leqslant t\leqslant T}$ satisfies Eq. (15), then $\mathbb{E}\left[\xi_{T}^{\boldsymbol{\nu}}X_{T}^{*}\right]\leqslant x_{0}$. Define the new probability measure $\hat{\mathbb{P}}$, with
We see that if $\{\boldsymbol{\nu}_{t}\}_{0\leqslant t\leqslant T}$ satisfies Eq. (15), $\hat{\mathbb{P}}$ is a well-defined probability measure and $\frac{\mathrm{d}\hat{\mathbb{P}}}{\mathrm{d}\mathbb{P}}$ is the Radon–Nikodym derivative. By Girsanov’s Theorem, the components of Brownian motion $\{\hat{\mathbf{W}}_{t}\}_{0\leqslant t\leqslant T}$ under $\hat{\mathbb{P}}$ is
$=\exp\left\{-\int_{0}^{T}\left(r_{s}+\frac{1}{2}\|\boldsymbol{\theta}_{s}\|_{2}^{2}\right)\,\mathrm{d}s+\int_{t}^{T}\boldsymbol{\theta}_{s}^{\intercal}\boldsymbol{\nu}_{s}\mathrm{d}s-\int_{0}^{T}\boldsymbol{\theta}_{s}^{\intercal}\,\mathrm{d}\mathbf{\hat{W}}_{s}\right\}$
This shows that $\xi_{T}$ has the same distribution under $\mathbb{P}$ and $\hat{\mathbb{P}}$. Note that $X_{T}^{*}=I\left(y^{*}\xi_{T}\right)$ is a function of $\xi_{T}$. Define $h(z):=zI\left(y^{*}z\right)$. Clearly, we have $\mathbb{E}\left[h(\xi_{T})\right]=\hat{\mathbb{E}}\left[h(\xi_{T})\right]$, where $\hat{\mathbb{E}}$ is the expectation under $\hat{\mathbb{P}}$. Combining with (61), it directly follows that
Lemma 1 and Proposition 5 show that if Eq. (15) holds, the optimality of $\{\boldsymbol{\pi}_{t}^{*}\}_{0\leqslant t\leqslant T}$ in Theorem 2 holds in incomplete markets.
5 Analysis on the Optimal Processes
In this section, we conduct an asymptotic analysis and a numerical analysis on the optimal wealth process and the optimal portfolio given in Section 4.1 to show some characteristics of these processes.
5.1 Asymptotic Analysis